Find the particular integral solution of d3y/dx3-4dy/dx=x+3cosx+e-2x

1
Expert's answer
2021-02-24T12:14:35-0500

CORRECTED SOLUTION

"\frac{d^3y}{dx^3}-4\frac{dy}{dx}=x+3 \cos x+e^{-2x}"

We have a linear non-homogeneous ODE with constant coefficients in its homogeneous part. The characteristic values are the roots of the algebraic equation:

"\lambda^3-4\lambda=0"

They are "\lambda_1=0" , "\lambda_2=2" and "\lambda_3=-2"

Each of the terms in the right-hand side x, 3cos x and e-2x are linear combinations of terms of the form "P(x)e^{\lambda x}" with P(X) is a polynomial. We may seek the partial solutions of the same form "Q(x)e^{\lambda x}" with degree of Q equal to degree of P, if the corresponding "\lambda" is non-characteristic, or greater by 1 (up to the multiplicity of "\lambda"), if else. Furthermore, as the order of the equation is odd (3), we are to replace cosine to sine.

So, let us seek a solution of the form "y=a_1x^2+a_2 \sin x+a_3xe^{-2x}". Then

"y'=2a_1x+a_2 \cos x-2a_3xe^{-2x}+a_3e^{-2x}"

"y''=2a_1-a_2 \sin x+4a_3xe^{-2x}-4a_3e^{-2x}"

"y'''=-a_2 \cos x-8a_3xe^{-2x}+12a_3e^{-2x}"

"y'''-4y'=-8a_1x-5a_2\cos x+8a_3e^{-2x}=x+3 \cos x+e^{-2x}"

Comparing the corresponding coefficients, we get:

a1=-1/8, a2=-3/5 and a3=1/8.


Answer. "y=-\frac{1}{8}x^2-\frac{3}{5} \sin x+\frac{1}{8}xe^{-2x}" .